when you reverse the digits in a certain two-digit number you decrease its value by 36. what is the number if the sum of its digits is 10.
i don't know how to set the equation up.
if you could that would be great! :)

Found 2 solutions by ewatrrr, sudhanshu_kmr:
Answer by ewatrrr(24785)


Hi
sum of its digits is 10
Let x and (10-x)represent the unit and tens digits of this certain two-digit number
Question states*** Note: [10*(10-x)+ x] the original number
10*x + (10-x) = [10*(10-x)+ x] - 36
Solving for x
9x + 10 = 100 - 9x - 36
18x = 54
x = 3, the units digit, 7 the ten digit, the Number is 73
CHECKING our Answer***
37 = 73 - 36

Answer by sudhanshu_kmr(1152)


let unit digit number is x , tenth digit is y...
x+y = 10 ..............(1)
original value of number is 10y + x
after reversing it is 10x + y
10y +x -( 10x +y ) = 36
=> 9y - 9x = 36
=> y - x = 4 (dividing by 4 both side)
now, add this equation to equation (1)
we find,
2y = 14
=> y = 7
and value of x = 3
so, original number is 7*10 + 3 = 73
after reversing it becomes 37